OpenStudy (el_arrow):

whats the derivative of ln(1/2)

OpenStudy (el_arrow):

i know you get -ln(2) i just dont understand how

OpenStudy (el_arrow):

like where does the negative sign come from?

OpenStudy (dumbcow):

is this a trick question? :) ln(1/2) is a constant so derivative = 0

OpenStudy (el_arrow):

no thats not it

OpenStudy (dumbcow):

well what do you mean by "derivative"

OpenStudy (el_arrow):

like on one of the problems i am working on it says that ln(1/2) equals -ln(2)

OpenStudy (dumbcow):

herea a log property \[\ln(x^n) = n*\ln x\] 1/2 can be written as 2^-1 \[\ln(1/2) = \ln(2^{-1}) = -\ln 2\]

OpenStudy (aum):

ln(A/B) = ln(A) - ln(B) ln(1/2) = ln(1) - ln(2) = 0 - ln(2) = -ln(2)

OpenStudy (el_arrow):

oh yes thats it

OpenStudy (el_arrow):

alright thanks

OpenStudy (dumbcow):

to avoid further confusion, please do not use the term derivative for these type of questions :)